OpenStudy (anonymous):

find the smallest perimeter and the dimensions for a rectangle with an area of 36 in^2

OpenStudy (anonymous):

let "b" be the base let "h" be the height area of a rectangle = 36 A= b*h 36= b*h Perimeter- want to minimize P=2b+2h substitute the area equation into the perimeter equation so you have perimeter in terms of only base or height then find the derivative of the perimeter- set it equal to zero- solve for b or h (depending on which variable you substituted)- plug b or h back into the perimeter equation to determine minimum perimeter